OpenStudy (anonymous):

Emily is considering two mutually exclusive financial options: (i) to deposit $4,000 in her bank saving account that pays 4.6% annual interest, or (ii) to purchase a $4,00 one-year guaranteed investment certificate with monthly interest rate of 0.1%. From the opportunity cost standpoint, by making the decision to DEPOSIT $4,000 in the bank account Emily will: A. lose $156 by the end of the year B. gain $156 by the end of the year C. lose $136 by the end of the year D. gain $136 by the end of the year E. make zero economic profit The answer is C. I just don't understand the method

OpenStudy (anonymous):

This is for my engineering economics class. This is a question from chapter two in the class, and the chapter deals with Interests and Interest Rates, Compound and Simple Interest, Effective and Nominal Interest Rates, Continuous Compounding, Cash Flow Diagrams and Equivalence. Possible equations:

OpenStudy (anonymous):

compound interest - F=P(1+i)^N simple interest - Is=PIN effective and nominal interest rates - if 'is'= (nominal interest rate 'r')/(compounding periods 'm') then F=P(1+'is')^m = P(1+ie) continuous compounding 'ie'=(e^r-1)

OpenStudy (anonymous):

@ganeshie8 @dan815 @bookworm00981 @Preetha @esshotwired @JoannaBlackwelder @jim_thompson5910 Let me know if any of you have some knowledge on the subject. If anyone could help it would be exceedingly appreciated!!!

OpenStudy (joannablackwelder):

For the bank account, you can use the simple interest formula since the interest is annual.

OpenStudy (joannablackwelder):

What do get get the amount of the interest to be for the bank account?

OpenStudy (anonymous):

Is= 4000(0.046)(1) = 184 P.S. I'm at hotel at the moment and the internet is working off and on, sorry if the reply take long.

OpenStudy (anonymous):

@JoannaBlackwelder I think that's right. Is = 184 so F = 4000+184= 4184

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

@anonymous_user Could you be of any help with this question? I don't know where @JoannaBlackwelder went.

OpenStudy (joannablackwelder):

Yes, good! Sorry, I was away for a while.

OpenStudy (joannablackwelder):

I think you would use compound interest for the investment.